OpenStudy (anonymous):

f(x)= sqrt(6x-6) and g(x)=6x^2-2 For each function h given below, find a formula for h(x) and the domain of h. Enter the domains using interval notation. (A) h(x) = (f o g)(x). h(x) = Domain = (B) h(x) = (g o f)(x). h(x) = Domain = (C) h(x) = (f o f)(x). h(x) = Domain = (D) h(x) = (g o g)(x). h(x) = Domain =

OpenStudy (anonymous):

Do you need help finding the composite function or the domain? To do the first problem you need to find the composite function (f o g)(x). That means find the function f(g(x)). So g(x)=6x^2-2. Substitute that in f(g(x)) and it's f(6x^2-2). Since f(x) = sqrt(6x-2), replace all x with 6x^2-2. You would get f(6x^2-2) = sqrt(6(6x^2-2)-2) Simplify the problem and it's sqrt(36^2-14) So f(6x^2-2)= sqrt(36^2-14). The domain would be all real numbers since parabolas never have a problem. Every x has a y.

OpenStudy (anonymous):

Then you just have go through the same process for the rest of the problems